OpenStudy (anonymous):

does any one knows how i can solve this: ( 4√3 - 6√2)(3√3 + 2√2) ?

OpenStudy (shamil98):

multiply using foil. (a+b)(c+d) = ac + ad + bc + bd

OpenStudy (anonymous):

36+8√6-18√6-24

OpenStudy (anonymous):

solve this u get the answer

OpenStudy (anonymous):

value of √6=2.45

OpenStudy (yttrium):

Hey @wolfine . It's like this. (a+b) (c+d) = ac+ad+bc +bd Hence \[(4\sqrt{3})(3\sqrt{3} )+ (4\sqrt{3})(2\sqrt{2}) + (-6\sqrt{2})(3\sqrt{3} )+(-6\sqrt{2})(2\sqrt{2})\]

OpenStudy (yttrium):

Can you evaluate that? If yes, then show it. :)

OpenStudy (anonymous):

as u get 12.5 something

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

@Yttrium

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

is 12-10radical 6

OpenStudy (anonymous):

12−106√

OpenStudy (anonymous):

12−10√6

OpenStudy (yttrium):

Yup! You're right! :)) I think it's better if you factor it out. It's factorable, right?

OpenStudy (anonymous):

thank you for answer my question guys